Find the solution to the equation
8^(4x+1)=16^(4x+2)

Answer by stanbon(75887)

Find the solution to the equation
8^(4x+1)=16^(4x+2)
(2^3)^(4x+1) = (2^4)^(4x+2)
2^(12x+3) = 2^(16x+8)
Equate the exponents to get:
12x+3= 16x+8
4x = -5
x = -5/4
============
Cheers,
Stan H.